A triangle has a perimeter of 24 yards. The shortest side is 4 yards shorter than the longest side, and the sum of the shortest and longest sides is equal to twice the length of the other side. What are the lengths of the three sides of this triangle?
Hi
I have the answer but I am getting stuck on the algebraic formula that would arrive at that answer. Can someone please lead me in the right direction?
Thanks
Matt

Answer by checkley77(12569)

x=z-4 0r x-z=4
x+z=2y
x+y+z=24
Replace (x+z) with 2y
2y+y=24
3y=24
y=24/3
y=8 ans.
x+z=2*8
x+z=16
x-z=4 add these 2 equations.
----------
2x=20
x=20/2
x=10 ans.
10-z=4
-z=4-10
-z=-6
z=6ans.
Proof:
8+10+6=24
24=24